Draw line 1 through (-4,0) and (0,6). What is the slope of any line parallel to line 1? Draw line 2 through the origin and parallel to line 1.
Answer by Alan3354(69443)

Draw line 1 through (-4,0) and (0,6). What is the slope of any line parallel to line 1?
Slope, m = diffy/diffx = (0-6)/(-4-0) = 3/2
All lines parallel have the same slope.
-----------------------------------
Draw line 2 through the origin and parallel to line 1.
You will have to do that. I did, and held it up to the screen, but you can't see it.